Find complex conjugate and the product of conjugates, for: sqrt(5)-4i
Answer by TimothyLamb(4379)

sqrt(5) - 4i
---
complex conjugate:
sqrt(5) + 4i
---
product of conjugates (p):
---
p = (sqrt(5) - 4i) * (sqrt(5) + 4i)
p = sqrt(5)*sqrt(5) + sqrt(5)*4i - sqrt(5)*4i - 4i*4i
p = 5 - 16ii
---
recall that i = sqrt(-1) and ii = -1
---
p = 5 - 16(-1)
p = 5 + 16
p = 21
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---